Circular No. 6177

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

MICROLENSING EVENT IN THE GALACTIC BULGE The OGLE microlensing search team (A. Udalski, M. Szymanski, J.

Kaluzny, M. Kubiak, W. Krzeminski, and M. Mateo; Warsaw University Observatory, Carnegie Institution of Washington, and University of Michigan) reports detection of another possible gravitational microlensing event in progress with the OGLE Early Warning System: “Observations collected with the 1-m Swope telescope at the Las Campanas Observatory indicate that the star (α = 18^{h}02^{m}07^{s}.59, δ = −30^{o}01′13′′.0, equinox 2000.0) had brightened by ∼ 0.7 mag from its normal brightness during 1992–1994 (I = 18.45, V − I = +1.55). The candidate seems to be on the rising branch of the microlensing light curve, well ahead of maximum. The event will very likely have a long timescale. We encourage follow-up photometric and spectroscopic observations that may provide valuable information about the microlensing object. A finding chart, light curve, and regularly updated photometry can be found in the OGLE archive (anonymous ftp

sirius.astrouw.edu.pl, directory /ogle/ews/1995/bw5).”

1991 JX M. Hicks and W. Grundy, Lunar and Planetary Laboratory, University

of Arizona, write: “CCD spectrograms (0.55–1.05 μm, resolution 1.5 nm) of the earth-approaching asteroid 1991 JX were taken at the Catalina Station 1.54-m telescope on May 29, 30, and 31. Analysis of the spectra shows that the relative reflectance of 1991 JX is in complete agreement with the spectral family of Vesta-like objects as identified by Binzel and Xu (1993, Science 260, 186). Our observations represent a total of 6.8 hr of integration. The composite spectrum shows a reddish slope, increasing by a factor of 12 percent over a range from 0.55 to 0.70 μm. The apparent pyroxene absorption band is centered at 0.92–0.94 μm, with the characteristic upturn at longer wavelengths.”

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Unfiltered CCD total magnitudeS reported by J. V. Scotti, Kitt Peak

(0.9-m Spacewatch telescope): May 23.20 UT, 21.4; 23.24, 21.0. The comet appeared very nearly stellar, with a coma diameter of 8′ on May 23.

1995 June 8 Daniel W. E. Green